OpenStudy (anonymous):

The area of a rectangular athletic field is represented by the expression 14z^4-42z^3+63z square meters. Write an algebraic expression to represent one possible set of dimensions (in the sense “length times width”) of the athletic field. Include correct units with your solution.

OpenStudy (anonymous):

Factor 14z^4-42z^3+63z

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

not sure how to do

OpenStudy (anonymous):

14z^4-42z^3+63z = z (14z^3 - 42z^2 + 63)

OpenStudy (anonymous):

So now, Area = length * breadth = (14z^3 - 42z^2 +63) * z

OpenStudy (anonymous):

that's the answer then?

OpenStudy (anonymous):

So, u can say length=14z^3 - 42 z^2 +63 meter And width =z meter

OpenStudy (anonymous):

that's all I need then?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)yw